# Differentiation, the Exception Not the Rule - Evidence for Full Miscibility in Sub-Neptune Interiors

### v1

@@ 5.1 Phase changes with P and T
Figure 9: Isothermal, isobaric projections of the Gibbs free energy of mixing surface in the MgSiO3-Fe-H2 system at six different temperatures and pressures along a nominal T-P path through our fiducial sub-Neptune planet. Concentrations are in mole fractions. The dark grey region is the region of spontaneouos exsolution into two phases and is defined by the spinodals (black curves). The blue curves are the binodals defining the coexisting compositions for two stable phases. Example pairs of coexisting compositions are shown by the tie lines between two binodes (blue points) on the binodals. The bulk composition of our reference sub-Neptune with 4 wtH2 is shown by the circled Neptune symbol. The bulk composition for Earth, including H in the metal core (Young et al., 2023) is shown for reference.
@@
The derived $P$ vs. radial position for this fiducial sub-Neptune is shown in Figure 8. A sequence of isothermal, isobaric phase diagrams starting near the surface of the magma ocean where a supercritical magma ocean and a hydrogen-rich envelope coexist, and ending deep in the interior, is shown in Figure 9. The location of each of the phase diagrams is indicated in Figure 8. Focusing on our fiducial sub-Neptune bulk composition, one sees a progression beginning with three phases being in equilibrium at $T=3500$ K and $P=4.0$ GPa (Figure 9 A), a hydrogen-bearing silicate, hydrogen-rich iron metal, and a more hydrogen-rich envelope. As pressure increases as we move deeper through the planet, the phase boundary separating the supercritical magma ocean from the envelope is breached, and our sub-Neptune interior is composed of hydrogen-rich liquid silicate and iron metal (Figure 9 B). Once pressures exceed about 20 GPa ($T\sim 5000$ K), the bulk composition corresponding to our fiducial sub-Neptune is composed of one phase (Figures 9 C and D). The transition at these conditions from two stable phases to one phase is at shallow depths in the planet (Figure 8). Deeper still, the island of two-phase stability (e.g., Marcilla et al., 2023) continues to shrink until at virtually any concentration of H2 greater than about $1$% by mass, the interior of the planets are composed of a single phase (Figure 9 E). After disappearing at higher pressures, the island of two-phase stability makes a reappearance at pressures greater than about 300 GPa along the fiducial isentrope (Figure 9 F), but the deepest interior of the fiducial sub-Neptune remains entirely miscible (Figure 9 F).
@@ 5.2 Sensitivity
Figure 10: Changes in the topology of the ternary phase diagram shown in Figure 9 D with variations in the interaction parameter $\rm L_{AC}=L_{CA}$ by $+20$% and $-20$%.
We can show, however, changes in phase diagram topologies associated with reasonable variations in the interaction parameters. Arguably, the most uncertain interaction parameter is that for the symmetrical mixing in the MgSiO3-Fe binary since it has been approximated using the MgO-Fe system. Insixiengmay & Stixrude (2025) report a fit uncertainty of 5% in this regular interaction parameter, but this is a measure of precision rather than potential systematic errors. We infer that the actual accuracy is likely lower by a factor of several, as is often the case in practice. Figure 10 shows the changes in topology in the ternary phase diagram shown in Figure 9 D with ±$\pm 20$% variations in interaction parameter $\rm L_{AC}=L_{CA}$, a spread that is four times the fit uncertainty. The changes are relatively minor in that the transition from two stable phases to complete miscibility changes by about 10 GPa in pressure along the isentrope, corresponding to a relatively small change in depth (Figure 8).
Similarly, $\sim 20$% variations in the other interaction parameters (and in many cases much greater variations) have relatively minor effects similar to those shown in Figure 10. Among the parameters comprising the ternary mixing model, the pressure dependence of the mixing in the Fe-H2 system has the most impact on the results. For separate silicate and Fe-rich metal phases to be stable at depth in our fiducial sub-Neptune with 2% H2 (e.g., at 100 GPa along the adiabat), the pressure-dependence of the interaction parameters in Equation 3 would have to be lower than the adopted value by at least 30%. Stabilization of the two phases results from expansion of the island of two-phase stability in the high-H2 portion of the phase diagrams (e.g., Figure 9 D). However, such a low pressure-dependence violates experimental and computational constraints on the distribution of H2 between silicate and Fe metal at relevant $T$ and $P$. As described above, $D_{\mathrm{H}}=[H]_{\mathrm{metal}}/[H]_{\mathrm{silicate}}$ is orders of magnitude greater than unity (e.g., Tagawa et al., 2021) at magma ocean conditions. This is not the case where $\partial L_{\rm AB}/\partial P$ is significantly less than the value of 9500 used in our model.
@@ 6 Discussion
$R/R_{\oplus}$ 3.485 2.923 2.756
$R_{\rm core}/R_{\oplus}$ 1.857 2.004 1.908
$M_{\rm core}/M_{\oplus}$ 5.960 5.947 5.951
@@
bulk planet $\rho$ (g $\rm cm^{-3}$) 0.781 1.324 1.580
mass fraction envelope 0.655% 0.873% 0.808%
$w_{\rm H_{2}}$ silicate core 1.33% 1.33% 2.12%
$w_{\rm H_{2}}$ envelope 100% 77.93% 72.26%
footnotetext: $T_{\rm s}=$ magma ocean surface temperature (K), $P_{\rm s}=$ magma ocean surface pressure (GPa), $w_{i}$ is the concentration of component $i$ by mass (where indicated).
Figure 11: Comparison of two models for a fiducial sub-Neptune. In both cases the planets have identical masses of $6M_{\oplus}$, equilibrium temperatures of 900K, and bulk H2 concentrations of 2% by mass. In panel A the planet is an “unequilibrated” structure composed of a pure Fe metalic liquid core overlain by an MgSiO3 melt with a specified concentration of H2 to match the fully-miscible case in panel B, in turn overlain by a pure H2-rich envelope. The silicate-envelope boundary is specified to be the same $T$ and $P$ as the binodal in panel B. The planet in panel B is composed of a supercritical magma ocean that is a fully miscible ternary mixture of MgSiO3, Fe, and H2. The surface of the magma ocean is defined by the MgSiO3-H2 binodal, where $T=3591$ K and $P=4.0$ GPa for the melt H2 concentration of 1.3% by mass. Note mass densities are given in kg $\rm m^{-3}$.
@@
where $\Delta R_{\rm A}$ and $\Delta R_{\rm B}$ are the thicknesses of the envelopes for the models in panels A and B in Figure 11. The first term on the right-hand side of Equation 9 is the ratio of molecular weights. The mean molecular weight of the envelope for the unequilibrated, pure H2 envelope is, by design, $2.3$ amu (e.g., H2 with an assumed He concentration) throughout, while the mean molecular weight of the envelope in the single-phase, fully-miscible interior model is $3.5$ amu near the base where much of the mass resides, transitioning upward to 2.3 amu at lower densities. The first term can thus be approximated as $3.5/2.3=1.52$. The second term can be evaluated using the radius and density data for the cores in Table 1, yielding a value of 0.857. The simplified prediction for the ratio $(R_{\rm p}-R_{\rm core})_{\rm A}/(R_{\rm p}-R_{\rm core})_{\rm B}$ from Equation 9 is $1.31$. The full model ratio is $1.77$ (Table 1), the difference being attributable to the effects of convective inhibition, a moist adiabat, and other factors not included in this simple analysis. These calculations illustrate that the greater density of the core in the unequilibrated case compared with the fully miscible core should decrease the radius of the planet, all else equal, but the difference in mean molecular weights in the lower atmospheres overwhelms this effect.
@@
In order to better isolate the effect of the core density on the bulk density of our fiducial $6M_{\oplus}$ planet, we compare the fully miscible interior model in Figure 11 B with a similar model in which the only difference is that there is a pure Fe, Earth-like metal core. The Fe metal and silicate are in 1:2 proportions. Again, the total planet mass and mass fraction of hydrogen are maintained. The result is shown in Figure 12. Here the effects on bulk density due to the characteristics of the differentiated core and the fully miscible core are evident. Presence of a pure Fe core interior to the silicate increases the overall core density of the planet from $4.08$ g $\rm cm^{-3}$ to $4.72$ g $\rm cm^{-3}$. The planet with the pure Fe central core has a smaller transit radius of $2.76$ $R_{\oplus}$ compared with $2.92$ $R_{\oplus}$ for the single-phase case, and thus a greater bulk density of 1.58 g $\rm cm^{-3}$ compared with the fully miscible case where the bulk density is $1.33$ g $\rm cm^{-3}$.
Although the difference in transit radii is impacted by core densities, here again the envelopes also contribute. The two models have different mass fractions of envelopes, and the envelopes differ in their mean molecular weights near the base. This is because the weight fraction of H2 in the supercritical silicate-H2 phase in the case of a pure Fe core is necessarily greater than that in the fully miscibile core case (again, to preserve the whole-planet H2 mass fraction), with values of $2.1$ and $1.3$%, respectively. As a consequence of the shape of the binodal curve, higher concentrations of H2 in the melt phase necessitate higher concentrations of heavy elements in the equilibrated gas, or envelope, phase (see Figure 2). The resulting maximum values for mean molecular weights just above the magma ocean surface are $5.2$ amu in the Fe-core case and $3.5$ amu in the fully-miscible core case. This results in a ratio $\Delta R_{\rm 11A}/\Delta R_{\rm 10B}$ of $<1$, with the precise value requiring integration over the depths of the envelopes. The important conclusion is that the envelope enhances the effect of different core densities on the average planet density in this case, rather than opposing the effects of core densities as in the previous case (Table 1).
@@
The result of this comparison of models in Figure 12 is that substituting a single, supercritical silicate-Fe-H2 phase for an Earth-like central Fe metal core and an overlying H-bearing silicate magma ocean has a marginal effect on the observable bulk density of the planet when the total mass fraction of H2 for the planet is the same, with a difference in radius of only 6%. This is despite the significant reduction in the density of the core as a whole (Table 1). What does change is the mass fraction of the hydrogen-rich envelope, and the concentration of hydrogen in the silicate portion of the core (Table 1). The sum of thermal and gravitational potential energies for the two planets are marginally different, by about 2%, with values of $-4.832\times 10^{33}$ and $-4.746\times 10^{33}$ Joules for the Fe-core and fully-miscibile cases, respectively.
@@ 7 Conclusions
We show that based on our present understanding of the phase equilibria of the MgSiO3-Fe-H2 system, the interior of sub-Neptunes are unlikely to be composed of a discrete silicate mantle and iron core during, and presumably after, their magma ocean phase of evolution. Discrete silicate and metal phases are expected to occur in only the very shallowest depths of these planets. The mixed-phase core is under-dense, with a compressed bulk density closer to 4 g $\rm cm^{-3}$ for a core mass of 2 $M_{\oplus}$ as compared with Earth’s compressed bulk density of 5.5 g $\rm cm^{-3}$. The boundary between this fully miscible core and the outer H2-rich envelopes corresponds to the phase change from a supercritical melt to coexisting melt and envelope. The base of the envelope is predicted to have significant fractions of heavy elements.
These results are important for motivating investigations into the equations of state for these core materials, as well as the chemistry of sub-Neptune cores. With our current assumptions of linear mixing of densities among the three components, and compressibilities similar to MgSiO3, the effect of the supercritical, one-phase cores on the bulk density of the planets is on the order of several percent. The dominant feature that controls the radius at a given mass remains the inventory of heavy elements in the atmosphere, as well as its thermal structure. But we have shown that miscibility is a mechanism for storing hydrogen throughout the cores of these planets that is distinct from extrapolations of solubilities of gases into silicate melt. The physical chemistry is much richer, and deserving of further exploration if we are to understand this important class of planets. With this in mind, it is tempting to extrapolate these results to the more massive Neptune-like planets as well; might ice giants contain within them supercritical mixtures dominated by Mg, Si, O, Fe and H? We are not aware of any data nor strong arguments against such a proposition.

### v4

@@ 5.1 Phase changes with P and T
Figure 10: Isothermal, isobaric projections of the Gibbs free energy of mixing surface in the MgSiO3-Fe-H2 system at six different temperatures and pressures along a nominal T-P path through our fiducial sub-Neptune planet. Concentrations are in mole fractions. The dark grey region is the region of spontaneouos exsolution into two phases and is defined by the spinodals (black curves). The blue curves are the binodals defining the coexisting compositions for two stable phases. Example pairs of coexisting compositions are shown by the tie lines between two binodes (blue points) on the binodals. The bulk composition of our reference sub-Neptune with 2 wt% H2 is shown by the circled Neptune symbol. The bulk composition for Earth, including H in the metal core (Young et al., 2023), as well as the bulk composition of Neptune, are also shown for reference.
For these calculations, we prescribe the mass of the planet, the total H2 abundance, the equilibrium temperature for the atmosphere, and the thermal state of the planet as expressed by the pressure of the binodal defining the outer boundary of the magma ocean, $P_{\rm S}$. The pressure of the binodal is a manifestation of the total energy of the planet. Solutions are obtained by iterating between the core and atmosphere structures as follows. As a starting point for the iteration, we assume all of the H2 resides in the core and calculate the surface temperature of the magma ocean, i.e. the binodal temperature where one phase becomes unstable relative to two stable phases, melt and envelope, at the prescribed concentration of H2. This temperature is the reduced temperature for the adiabat in the core (i.e., the surface $T_{\rm S}$ for given $P_{\rm S}$). The shape of the binodal determines the composition of the atmosphere in equilibrium with the melt at $T_{\rm S}$ and $P_{\rm S}$ (Figure 2). From the pressure at the base of the atmosphere (the pressure of the binodal) and its composition, we calculate an atmosphere model, including its mass. The lower boundary conditions for the atmosphere are set by the magma-envelope binodal, and the upper boundary is set by the equilibrium temperature. We then subtract the mass of hydrogen in this atmosphere from the core, and revise the core calculation, repeating for several iterations until the planet structure converges. The derived $P$ vs. radial position for this fiducial sub-Neptune is shown in Figure 9.
@@
Using our ternary Gibbs free energy model, we constructed a sequence of isothermal, isobaric phase diagrams starting near the surface of the magma ocean where a supercritical magma ocean and a hydrogen-rich envelope coexist, and ending deep in the interior. The results are shown in Figure 10. The location of each of the phase diagrams in our model sub-Neptune is indicated in Figure 9. Focusing on our fiducial sub-Neptune bulk composition, one sees a progression beginning with three phases being in equilibrium at $T=3500$ K and $P=3.8$ GPa (Figure 10 A), a hydrogen-bearing silicate, hydrogen-rich iron metal, and a more hydrogen-rich envelope. As pressure increases deeper through the planet, the phase boundary separating the supercritical magma ocean from the envelope is breached, and our sub-Neptune interior is composed of hydrogen-rich liquid silicate and iron metal (Figure 10 B). Once pressures exceed about 20 GPa ($T\sim 5000$ K), all compositions with greater than approximately 1% by mass H2, including the bulk composition corresponding to our fiducial sub-Neptune, are composed of one phase (Figures 10 C and D). The transition at these conditions from two stable phases to one phase is at shallow depths in the planet, at $R/R_{\rm c}>0.95$ (Figure 9). Deeper still, the interiors of all planets composed of silicate, iron, and about 1% or greater hydrogen by mass remain as a single phase (Figures 10 E and F). The interiors of sub-Neptunes will have lower mass fractions of hydrogen than the bulk planet values because of the H2 comprising the envelopes. Our models show that for a bulk of 2% by mass H2, the interiors of sub-Neptunes similar to our fiducial case are expected to have roughly 1.31.3% by mass H2 (Table 1, see discussion below), placing them above the $\approx 1$% threshold separating the two-phase region from the one-phase region at the highest pressures in Figure 10.
This is the main result of these calculations; the interiors of planets with typical proportions of molten rock and iron, when mixed with H2 at roughly weight percent levels, form a single phase, and do not differentiate into a metal core and silicate mantle. Weight percent levels of total hydrogen are expected for sub-Neptunes, as described in the Introduction.
@@ 5.2 Sensitivity
Figure 11: Changes in the topology of the ternary phase diagram shown in Figure 10 D with variations in the interaction parameter $\rm L_{AC}=L_{CA}$ by +20+20% and 20-20%.
@@
We can show, however, changes in phase diagram topologies associated with reasonable variations in the interaction parameters. Arguably, the most uncertain interaction parameter is that for the symmetrical mixing in the MgSiO3-Fe binary since it has been approximated using the MgO-Fe system. Insixiengmay & Stixrude (2025) report a fit uncertainty of 5% in this regular interaction parameter, but this is a measure of precision rather than potential systematic errors. We infer that the actual accuracy is likely lower by a factor of several, as is often the case in practice. Figure 11 shows the changes in topology in the ternary phase diagram shown in Figure 10 D with $\pm 20$% variations in interaction parameter $\rm L_{AC}=L_{CA}$, a spread that is four times the fit uncertainty. The changes are relatively minor. The most significant change is that with a greater value for $\rm L_{AC}$, an island of two-phase stability appears near the composition of our fiducial sub-Neptune. Compositions that fall within this island will be composed of two phases, with their relative proportions given by the lever rule. However, the two phases are much more similar in composition to one another than silicate mantle and iron. One can see this by comparing the relatively short tie lines in the two-phase island to longer tie lines connecting silicate and metal in the two-phase field occupied by Earth (Figure 11). The appearance and disappearance of this two-phase island is a persistent manifestation of significantly altering the Gibbs free energy surface in these calculations.
Similarly, $\sim 20$% variations in the other interaction parameters (and in many cases much greater variations) have relatively minor effects similar to those shown in Figure 11. Among the parameters comprising the ternary mixing model, the pressure dependence of the mixing in the Fe-H2 system has the most impact on the results. For separate silicate and Fe-rich metal phases to be stable at depth in our fiducial sub-Neptune with 2% H2 (e.g., at 100 GPa along the adiabat), the pressure-dependence of the interaction parameters in Equation 4 would have to be lower than the adopted value by at least 30%. Stabilization of the two phases results from emergence of the island of two-phase stability in the high-H2 portion of the phase diagrams (e.g., Figures 8 and 11). However, such a low pressure-dependence violates experimental and computational constraints on the relative distributions of H2 between silicate and Fe metal at relevant $T$ and $P$, as described above.
@@ 6 Discussion
The complete miscibility of silicate, iron metal, and hydrogen in sub-Neptunes leads to a relatively under-dense core as compared with Earth-like layered models. As an illustration, we constructed models that can be used to compare the structures of sub-Neptunes composed of a completely miscible phase in their deep interiors with the more classical “unequilibrated” interiors composed of an inner liquid Fe iron core and an outer liquid silicate mantle. As a predicate, we accept the premise that the interiors of sub-Neptunes should inherit significant mass fractions of hydrogen as a result of the accretion process, and that ingassing of their hydrogen-rich atmospheres after final assembly of the planets may not be the only, or even the primary, mechanism for conveying hydrogen to their interiors. Specifically, we assume that planets form from planetary embryos that accreted hydrogen-rich envelopes that in turn mixed with their molten interiors. The likelihood that primary hydrogen-rich atmospheres intermingled with magma oceans during planet formation has been studied in the context of Earth’s formation (Mizuno et al., 1980; Harper & Jacobsen, 1996; Jaupart et al., 2017; Olson & Sharp, 2018; Young et al., 2023) and in the context of sub-Neptune formation and evolution (Chachan & Stevenson, 2018; Kite et al., 2016, 2019; Schlichting & Young, 2022). The mass fraction of H2 that could be accreted from a surrounding protoplanetary disk by a $0.5M_{\oplus}$ embryo is of order 3% (Sasaki & Nakazawa, 1990; Lee & Chiang, 2015; Ginzburg et al., 2016), with more massive bodies accreting greater mass fractions of hydrogen. The fact that sub-Neptunes do not typically have the maximum mass fractions of H2 permitted by the accretion process alone speaks to the importance of various loss mechanisms during and after accretion (e.g., Biersteker & Schlichting, 2019; Owen et al., 2020).
@@
$R/R_{\oplus}$ 3.485 2.895 2.618
$R_{\rm core}/R_{\oplus}$ 1.857 1.988 1.827
$M_{\rm core}/M_{\oplus}$ 5.960 5.949 5.962
@@
bulk planet $\rho$ (g $\rm cm^{-3}$) 0.781 1.363 1.844
mass fraction envelope 0.655% 0.850% 0.808%
$w_{\rm H_{2}}$ silicate core 1.33% 1.34% 1.52%
$w_{\rm H_{2}}$ envelope 100% 77.78% 76.55%
Figure 12: Comparison of two models for a fiducial sub-Neptune. In both cases the planets have identical masses of $6M_{\oplus}$, equilibrium temperatures of 900K, and bulk H2 concentrations of 2% by mass. In panel A the planet is an “unequilibrated” structure composed of a pure Fe metalic liquid core overlain by an MgSiO3 melt with a specified concentration of H2 to match the fully-miscible case in panel B, in turn overlain by a pure H2-rich envelope. The silicate-envelope boundary is specified to be the same $T$ and $P$ as the binodal in panel B. The planet in panel B is composed of a supercritical magma ocean that is a fully miscible ternary mixture of MgSiO3, Fe, and H2. The surface of the magma ocean is defined by the MgSiO3-H2 binodal, where $T=3594$ K and $P=4.0$ GPa for the melt H2 concentration of 1.3% by mass. Note mass densities are given in kg $\rm m^{-3}$.
@@
where $\Delta R_{\rm A}$ and $\Delta R_{\rm B}$ are the thicknesses of the envelopes for the models in panels A and B in Figure 12. The first term on the right-hand side of Equation 10 is the ratio of molecular weights. The mean molecular weight of the envelope for the unequilibrated, pure H2 envelope is, by design, $2.3$ amu (e.g., H2 with an assumed He concentration) throughout, while the mean molecular weight of the envelope in the single-phase, fully-miscible interior model is $3.5$ amu near the base where much of the mass resides, transitioning upward to 2.3 amu at lower densities. The first term can thus be approximated as $3.5/2.3=1.52$. The second term can be evaluated using the radius and density data for the cores in Table 1, yielding a value of 0.871. The simplified prediction for the ratio $(R_{\rm p}-R_{\rm core})_{\rm A}/(R_{\rm p}-R_{\rm core})_{\rm B}$ from Equation 10 is $1.33$. The full model ratio is $1.79$ (Table 1), the difference being attributable to the effects of convective inhibition, a moist adiabat, and other factors not included in this simple analysis. These calculations illustrate that the greater density of the core in the unequilibrated case compared with the fully miscible core should decrease the radius of the planet, all else equal, but the difference in mean molecular weights in the lower atmospheres overwhelms this effect.
@@
In order to better isolate the effect of the core density on the bulk density of our fiducial $6M_{\oplus}$ planet, we compare the fully miscible interior model in Figure 12 B with a similar model in which the only difference is that there is a pure Fe, Earth-like metal core. The Fe metal and silicate are in 1:2 proportions. Again, the total planet mass and mass fraction of hydrogen are maintained. The result is shown in Figure 13. Here the effects on bulk density due to the characteristics of the differentiated core and the fully miscible core are evident. Presence of a pure Fe core interior to the silicate increases the overall core density of the planet from $4.18$ g $\rm cm^{-3}$ to $5.39$ g $\rm cm^{-3}$. The planet with the pure Fe central core has a smaller transit radius by about 10% compared with the fully miscible model, with values of $2.62$ $R_{\oplus}$ and $2.90$ $R_{\oplus}$ for the two bodies, respectively. The pure Fe central core case results in a greater bulk density for the planet of 1.84 g $\rm cm^{-3}$ compared with the fully miscible case where the bulk density is $1.36$ g $\rm cm^{-3}$, an increase of 35%.
Although the difference in transit radii is impacted by core densities, here again the envelopes also contribute. However,the important conclusion is that the envelopes enhance the effects of different core densities on the average planet density in this case, rather than opposing the effects of core densities as in the previous case (Table 1).
@@
The result of this comparison of models in Figure 13 is that substituting a single, supercritical silicate-Fe-H2 phase for an Earth-like central Fe metal core and an overlying H-bearing silicate magma ocean affects the observable bulk density of the planet when the total mass fraction of H2 for the planet is the same, with a difference in radius of order 10%. An even more significant change is the differences in the central core temperatures. The central temperature of the Fe-core case is about double that of the fully miscible case. The sum of thermal and gravitational potential energies for the two planets are therefore also different, by about 30%, with values of $3.872\times 10^{33}$ and $4.995\times 10^{33}$ Joules for the fully-miscibile and Fe-core cases, respectively. While they are of same mass and the same total H2 mass fraction, these planets would have arrived at their respective configurations along very different evolutionary paths.
@@ 7 Conclusions
We show that based on our present understanding of the phase equilibria of the MgSiO3-Fe-H2 system, the interior of sub-Neptunes are unlikely to be composed of a discrete silicate mantle and iron core during, and presumably after, their magma ocean phase of evolution. Discrete silicate and metal phases are expected to occur in only the very shallowest depths of these planets. The mixed-phase core is under-dense, with a compressed bulk density closer to 4.2 g $\rm cm^{-3}$ for a core mass of 2 $M_{\oplus}$ as compared with Earth-like compressed bulk densities of 5.4 g $\rm cm^{-3}$ that obtain for analogous planets with discrete metal cores. The boundary between this fully miscible core and the outer H2-rich envelopes corresponds to the phase change from a supercritical melt to coexisting melt and envelope. The base of the envelope is predicted to have significant fractions of heavy elements.
These results are important for motivating investigations into the equations of state for these core materials, as well as the chemistry of sub-Neptune cores. With our current assumptions of linear mixing of densities among the three components, and compressibilities similar to MgSiO3, the effect of the supercritical, one-phase cores on the bulk density of the planets is on the order of 30%. We have shown that miscibility is a mechanism for storing hydrogen throughout the cores of these planets that is distinct from extrapolations of solubilities of gases into silicate melt. The physical chemistry is much richer, and deserving of further exploration if we are to understand this important class of planets. With this in mind, it is tempting to extrapolate these results to the more massive Neptune-like planets as well; might ice giants contain within them supercritical mixtures dominated by Mg, Si, O, Fe and H? We are not aware of any data nor strong arguments against such a proposition.